OpenStudy (anonymous):

if one number is 6 more than another and the difference of their squares is 144, what are the numbers? please explain. thx!

OpenStudy (anonymous):

ok thx :)

OpenStudy (wolf1728):

num = num +6 Now let's square both numbers num^2 and num^2 + 12*num + 36 The difference equals 144 -12*num - 36 = 144 12*num + 36 = -144 12*num = 180 num = 15 other num = 9 CHECK 15^2 - 9^2 = 225 -81 = 144

OpenStudy (anonymous):

ok thank you :) :) :)